# CAS-ViT: Convolutional Additive Self-attention Vision Transformers for Efficient Mobile Applications

### v1

@@ Results
We conduct semantic segmentation experiments on ADE20K (Zhou et al. 2019) dataset. The proposed model is evaluated with Semantic FPN (Kirillov et al. 2019) with the normalization layers frozen, taken as backbone and loaded the pre-trained weights of ImageNet-1K classification. Following the common practice (Yu et al. 2022), the network is trained for 40K iterations on batch size of 32 and AdamW optimizer. The learning rate is initialized as 210-4 and decays in polynomial schedule with power of 0.9. Images are resized and cropped to 512512 during training. The model is implemented on MMSegmentation 333https://github.com/open-mmlab/mmsegmentation codebase.
Results
@@
The qualitative and quantitative analysis of segmentation results on the ADE20K dataset, as illustrated in Table 2, reveals the superior balance between computational efficiency and segmentation accuracy achieved by our models (CAS-ViT-XS, CAS-ViT-S, and CAS-ViT-M). Despite their relatively lower parameter counts and Flops, our models demonstrate competitive or superior mean Intersection over Union (mIoU) percentages, indicative of their effectiveness in semantic segmentation tasks. Specifically, CAS-ViT-M outperforms more computationally intensive models with an mIoU of 43.6%, underscoring its architectural efficiency and the efficacy of its design in capturing detailed semantic information. This balance of high efficiency and accuracy across our model variants highlights their suitability for diverse segmentation applications, especially in resource-constrained scenarios, establishing them as compelling options for efficient and effective semantic segmentation on ADE20K (Zhou et al. 2019) dataset.
Visualization
Fig. 4 presents a comparison of the heatmap visualizations carried out at the last layer of the networks, where the three methods share a similar parameter number. It can be observed that our method is able to accurately localize the region of interest to the critical part. Meanwhile in comparison, we obtain a larger receptive field, which is beneficial to overall performance, especially for subsequent dense prediction tasks. More visualizations on object detection and instance segmentation please refer to Appendix.
Ablation Study
@@
We demonstrate the effectiveness of the proposed method by ablation study. As shown in Table 4(#1-#3), we replace CATM with Pooling (Yu et al. 2022) and Window-MSA (Dosovitskiy et al. 2020), respectively. Pooling without information interaction -1.72%, while making interaction through W-MSA -0.02%. It verifies the feasibility of CATM’s information interaction via Sigmoid-activated attention. Table 4(#4-#5) explored the effect of lacking a dimension interaction on the results, with the absence of spatial and channel decreasing by 0.45% and 0.21%. (#6-#7) replace the context mapping function in the Query and Key branches with different ones, and only retaining the full interaction in #7 achieves similar results.
Limitation and Future Work
@@
Conclusion
In this paper, we proposed a convolutional additive self-attention network referred to as CAS-ViT. Firstly we argue that the key to enabling token mixer in ViT to work efficiently is multiply information interaction including spatial and channel domains. Subsequently following this paradigm, we innovatively designed an additive similarity function and implement it simply by Sigmoid-activated attention, which effectively avoids complex operations such as matrix multiplication and Softmax in vanilla ViTs. We constructed a family of lightweight models and verified their superior performance on tasks: image classification, object detection, instance/semantic segmentation. Meanwhile, we have deployed the network to ONNX and Iphone, and the experiments demonstrate that CAS-ViT maintains high accuracy while facilitating efficient deployment and inference on mobile devices.

### v2

@@ IV-C1 Implementation Details
We conduct semantic segmentation experiments on ADE20K [67] dataset which contains 25K images covering 150 fine-grained categories. The proposed model is evaluated with Semantic FPN [68] with the normalization layers frozen, taken as backbone, and loaded the pre-trained weights of ImageNet-1K classification. Following the common practice [7], the network is trained for 40K iterations on batch size of 32 and AdamW optimizer. The learning rate is initialized as 210-4 and decays in a polynomial schedule with a power of 0.9. Images are resized and cropped to 512512 during training. The model is implemented on MMSegmentation 333https://github.com/open-mmlab/mmsegmentation codebase.
Backbone | Par.(M) | Flops(G) | mIoU(%) |
CAS-ViT-XS | 6.9 | 24.2 | 37.1 |
EdgeViT-XXS | 7.9 | 24.4 | 37.9 |
EfficientFormer-L1 | 15.6 | 28.2 | 38.9 |
PVT-S | 28.2 | 44.5 | 39.8 |
PoolFormer-S24 | 25.1 | 39.3 | 40.3 |
CAS-ViT-S | 9.4 | 26.1 | 41.3 |
SwiftFormer-L1 | 15.5 | 29.7 | 41.4 |
PoolFormer-S36 | 34.6 | 47.5 | 42.0 |
EfficientFormerV2-S2 | 16.3 | 27.8 | 42.4 |
PVTv2-B1 | 17.8 | 34.2 | 42.5 |
CAS-ViT-M | 15.8 | 31.3 | 43.6 |
PVT-L | 65.1 | 79.6 | 42.1 |
PoolFormer-M48 | 77.1 | 82.1 | 42.7 |
EfficientFormer-L3 | 34.7 | 42.0 | 43.5 |
SwiftFormer-L3 | 31.7 | 42.5 | 43.9 |
XCiT-S24 P16 | 51.8 | - | 44.6 |
CAS-ViT-T | 25.0 | 40.39 | 45.0 |
IV-C2 Results
@@
The quantitative analysis of segmentation results on ADE20K dataset, as illustrated in Table VI, reveals the superior balance between computational efficiency and segmentation accuracy achieved by our models. Despite their relatively lower parameters and Flops, our models demonstrate competitive or superior mean Intersection over Union (mIoU) percentages, indicative of their effectiveness in semantic segmentation tasks. Specifically, our M and T models achieve at 45.0% and 43.6% mIoU, outperforming recent work such as PVT [17], EfficientFormer [54], SwiftFormer [20], etc., with fewer number of parameters. The superior performance underscores its architectural efficiency and the efficacy of its design in capturing detailed semantic information. This balance of high efficiency and accuracy across our model variants highlights their suitability for diverse segmentation applications, especially in resource-constrained scenarios, establishing them as compelling options for efficient and effective semantic segmentation on ADE20K [67] dataset.
IV-D Visualization
IV-D1 Heatmap
Fig. 4 presents a comparison of the heatmap visualizations carried out at the last layer of the networks, where the three methods share a similar parameter number. It can be observed that our approach is able to localize the key information in the graph very accurately, with attention to objects across animals, buildings, cars, and so on. Meanwhile, in comparison with two H-MSA-based methods, PoolFormer [7] and SwiftFormer [20], our method is able to extract useful information over a wider range. This also confirms our analysis on receptive fields, where the information association of feature maps is achieved in the form of convolution and does not lose the global perceptual ability to a great extent. In addition, a stronger perceptual capability also benefits the overall performance, especially for intensive prediction tasks such as semantic segmentation.
IV-D2 Detection and Segmentation Results
Fig. 5 visualize the results of object detection/instance segmentation and semantic segmentation. Our method has the following advantages compared to PoolFormer: 1. There are fewer false detections, such as the skateboard and the cat in Fig. 5; 2. There is a significant improvement in detecting objects with edges and shaded objects, e.g., the bus; 3. For large objects, such as elephants in Fig. 5, it is able to retrieve and obtain the complete bounding box in the global range.
IV-E Ablation Study
Similarity Function | Param.(M) | Flops(G) | Top-1(%) | |
1 | Base | 12.4 | 1.88 | 80.43 |
2 | w/o spatial | 12.4 | 1.86 | 79.98 |
3 | w/o channel | 11.0 | 1.88 | 80.22 |
4 | , | 11.7 | 1.87 | 80.04 |
5 | 12.4 | 1.88 | 80.40 |
We demonstrate the effectiveness of the proposed method by ablation study. Firstly, we verify the effectiveness of CATM by replacing the token mixers. As shown in Fig. 6, we replace the CATM proposed in this paper with Pooling from ref [7] and Window-MSA. Pooling brings the benefit of Flops decrease and also a significant performance loss. While comparing with W-MSA, CATM is able to have higher accuracy while retaining lower Flops. It verifies the feasibility of CATM’s information interaction via Sigmoid-activated attention.
@@ IV-E Ablation Study
Then, as shown in Table VII, we verify the arrangement of multiple information interactions in CATM to ensure that the network performance is sufficiently exploited. Table VII(#2-#3) explored the effect of lacking a dimension interaction on the results, with the absence of spatial and channel decreasing by 0.45% and 0.21%. (#4-#5) replace the context mapping function in the Query and Key branches with different ones, and only retaining the full interaction in #7 achieves similar results. It can be concluded from this that multiple information interactions are necessary for the network to fully utilize its performance, while the arrangement order of spatial and channel domains is not particularly required.
V Limitation and Future Work
@@
VI Conclusion
In this paper, we proposed a convolutional additive self-attention network referred to as CAS-ViT. Firstly we argue that the key to enabling token mixer in ViT to work efficiently is multiply information interaction including spatial and channel domains. Subsequently following this paradigm, and for the network to be more readily deployed at the various edge devices, we innovatively designed an additive similarity function and implement it simply by Sigmoid-activated attention, which effectively avoids complex operations such as matrix multiplication and Softmax in vanilla ViTs. We constructed a family of lightweight models and verified their superior performance on tasks: image classification, object detection, instance/semantic segmentation. Meanwhile, we have deployed the network to ONNX and Iphone, and the experiments demonstrate that CAS-ViT maintains high accuracy while facilitating efficient deployment and inference on mobile devices.